Circular No. 9114

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A. cbatiau@eps.harvard.edu ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

U SCORPII H. L. Worters, South African Astronomical Observatory (SAAO); S.

P. S. Eyres and M. T. Rushton, University of Central Lancashire; and B. Schaefer, Louisiana State University, report that the resumption of optical flickering in the recurrent nova U Sco was detected on Feb. 5.174–5.222 UT, only eight days after the most recent outburst (Jan. 28.4743 UT; cf. IAUC 9111). Flickering was not apparent in any of the observations on each night from Jan. 29 to Feb. 3 with the SAAO 1.9-m telescope and CCD camera. The amplitude of flickering is around 0.2 mag over the course of an hour, consistent with accretion resuming in the system. This means that the ejecta are now becoming optically thin, and an accretion disk has already been re-established, placing a lower limit of 7 days on this event. Further observations are urged to follow development of this rapidly changing recurrent nova, including characterizing eclipses that should now be visible again.

COMET 29P/SCHWASSMANN-WACHMANN Regarding the outburst of this comet that was noted on CBET 2160,

D. T. Durig, University of the South, Sewanee, TN, reports that astrometric CCD images taken with a 30-cm reflector by Durig with six other observers show a brightening from mag 15–16 on Jan. 27.2 UT to mag 12.1–12.3 on Feb. 4.14–4.17. From astrometry reported to the Minor Planet Center, the last pre-outburst CCD magnitude (16.1) appears to be from Feb. 1.95 (J. Camarasa, Sabadell, Spain).

COMETS C/2009 N2, C/2009 N3, C/2009 O5, C/2009 P3 (SOHO) Additional presumed comets have been found on SOHO website images

— Kreutz sungrazers except for C/2009 N3 (non-group). C/2009 N2 was small and stellar in appearance (mag ∼ 7–7.5) in C3 images, but small and slightly diffuse in C2 images. C/2009 N3 was stellar in appearance and brighter (mag ∼ 6.5–7 at ∼ July 5.04 UT). C/2009 O5 was also stellar in appearance (mag ∼ 6.5). C/2009 P3 was slightly elongated (mag ∼ 5.5–6).

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 N2 July 4.179 6^{h}42.^{m}5 +20^{o}51′ C3/2 MK 2009-P20
C/2009 N3  4.896 6 47.8 +22 59 C2  BZ 2009-P20
C/2009 O5  29.154 8 16.1 +17 01 C3  BZ 2009-W09
C/2009 P3 Aug. 1.196 8 23.5 +15 28 C3  MK 2009-W09
2010 February 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green